# GOAT-Bench: Safety Insights to Large Multimodal Models through Meme-Based Social Abuse

### v2

@@ Abstract
The exponential growth of social media has profoundly transformed how information is created, disseminated, and absorbed, exceeding any precedent in the digital age. Regrettably, this explosion has also spawned a significant increase in the online abuse of memes. Evaluating the negative impact of memes is notably challenging, owing to their often subtle and implicit meanings, which are not directly conveyed through the overt text and imagery. In light of this, large multimodal models (LMMs) have emerged as a focal point of interest due to their remarkable capabilities in handling diverse multimodal tasks. In response to this development, our paper aims to thoroughly examine the capacity of various LMMs (e.g. GPT-4V) to discern and respond to the nuanced aspects of social abuse manifested in memes. We introduce the comprehensive meme benchmark, GOAT-Bench, comprising over 6K varied memes encapsulating themes such as implicit hate speech, sexism, and cyberbullying, etc. Utilizing GOAT-Bench, we delve into the ability of LMMs to accurately assess hatefulness, misogyny, offensiveness, sarcasm, and harmful content. Our extensive experiments across a range of LMMs reveal that current models still exhibit a deficiency in safety awareness, showing insensitivity to various forms of implicit abuse. We posit that this shortfall represents a critical impediment to the realization of safe artificial intelligence. The GOAT-Bench and accompanying resources are publicly accessible at https://goatlmm.github.io/, contributing to ongoing research in this vital field.
@@ 1 Introduction
Remarkable progress has been made in the field of natural language processing (NLP), marked by the emergence of large language models (LLMs) equipped with billions of parameters and pre-trained on extensive data (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a, b). Through instruction fine-tuning and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), cutting-edge LLMs, like ChatGPT and GPT-4 (OpenAI, 2023), have demonstrated exceptional proficiency in comprehending and executing human instructions, even without tuning the parameters to offer powerful zero-shot performance (Wei et al., 2021; Kojima et al., 2022; Lin et al., 2023b; Wei et al., 2023). In a related vein, efforts have been directed towards equipping large multimodal models (LMMs) (Team et al., 2023; OpenAI, 2023; Zhu et al., 2023; Liu et al., 2023a; Gong et al., 2023; Dai et al., 2023) with similar capabilities, extending their proficiency to a range of multimodal tasks (Hudson and Manning, 2019; Lu et al., 2022; Fu et al., 2023). This dual advancement in both unimodal and multimodal models marks a significant stride in the field of artificial intelligence, paving the way for more integrated and versatile applications.
@@
Recent literature (Huang et al., 2023) has highlighted that LLMs can achieve remarkable performance in addressing the social abuse inherent in natural languages (An et al., 2021). This area of application is notably more ‘subjective’ compared to other NLP tasks and presents significant challenges, even for human evaluators, who often demonstrate lower agreement rates. In addition to language expression, the proliferation of social media has led to the rise of a distinctive multimodal phenomenon: memes (Zenner and Geeraerts, 2018). A typical meme consists of an image coupled with a brief textual component (Pramanick et al., 2021a; Lin et al., 2023a). Their ease of sharing enables memes to rapidly disseminate across various online platforms. While commonly perceived with humor, memes are increasingly being co-opted for online abuse, utilizing a strategic interplay of text and imagery that often relates to current political and socio-cultural events (Sharma et al., 2022). This evolving use of memes highlights their potential impact on the digital landscape, especially in the context of contemporary discourse. The challenge in detecting social abuse in memes lies in the subtlety of their content, where the underlying meanings are not immediately apparent in the text and images of the memes (Pramanick et al., 2021b). This task demands extensive commonsense knowledge and reasoning skills, built upon a precise interpretation of both the textual and visual elements of memes.
@@
However, in contrast to LLMs, the abilities and constraints of LMMs in handling relatively subjective and nuanced tasks, particularly in identifying online abusive memes within multimodal inputs, are yet to be fully explored (Awadalla et al., 2023; Chen et al., 2023). These tasks are integral to understanding social phenomena, demanding nuanced social judgment and decision-making skills. Given that LMMs are trained on extensive and diverse image-text corpora, showcasing impressive generalization skills (Liu et al., 2023a), it becomes imperative to assess both the potential strengths and challenges they may face in addressing meme-based social abuse. This line of inquiry is especially pertinent to safety insights, focusing on how LMMs discern and process the intricate combination of visual and textual elements in a manner that is both accurate and socially conscientious so that its output does not cause harm or discomfort.
@@
To this end, we introduce the GOAT-Bench111 Goat is often depicted as evil in memes due to its historical and mythological associations with devilish figures, reflected in its unique features like horns and eyes., a comprehensive and specialized dataset designed to evaluate large multimodal models through meme-based multimodal social abuse. GOAT-Bench comprises over 6K diverse memes, encompassing a range of themes including hate speech and offensive content. Our focus is to assess the ability of LMMs to accurately identify online abuse, specifically in terms of hatefulness, misogyny, offensiveness, sarcasm, and harmfulness. We meticulously control for the granularity of each specific meme task to facilitate a detailed analysis. Furthermore, we extend our evaluation to assess the effectiveness of thought chains in discerning the underlying implications of memes for deducing their potential threat to safety.
@@
Leveraging the GOAT-Bench, we assess the capability of various LMMs as exemplified in Figure 1, to detect meme-based social abuse through complex reasoning. We conceive of GOAT-Bench as a comprehensive suite of benchmarks, with subsets tailored to evaluate specific model abilities and analyze the key strengths and weaknesses of LMMs. For instance, in tasks focused on sexism, such as misogyny, we observe significant disparities among LMMs; in broader tasks like harmfulness, which demand advanced background knowledge and reasoning, the performance of LMMs tends to be more concentrated and generally modest. Our aim with GOAT-Bench is to enable researchers to gain a multi-dimensional understanding of their models’ abilities in addressing meme-based social abuse and to contribute to the advancement of safety insights in LMMs.
@@ 2.1 Tasks
Task III: Offensiveness. Offensiveness spreads an idea or emotion that intends to damage the social identity of the target person, community, or lower their prestige (Suryawanshi et al., 2020). In contrast to hateful memes, which target specific groups with expressions of animosity or discrimination, offensive memes encompass a wider spectrum of language or imagery that has the potential to offend, hurt, or provoke anger among individuals. They may not be directed towards a particular group and can affect individuals in a general sense (Drakett et al., 2018). In light of this understanding, it becomes imperative to investigate the efficacy of LMMs in scrutinizing online multimodal offensive meme content, known for its intent to provoke distress or embarrassment through its use of rudeness. To this end, we organized the offensive memes by utilizing the MultiOFF dataset (Suryawanshi et al., 2020) related to the 2016 United States presidential election, which consists of abusive and discriminatory messages against a person or minority group.
@@
Term Scope Sources Label Distribution Total
@@
Misogyny Targets women with sexism or hate, involving shaming, stereotyping, objectification, or violence. MAMI Misogynistic 500 1,000
@@
Sarcasm Feature irony and humor to subtly express dissatisfaction or mock, often through incongruous text and imagery. MSD Sarcastic 910 1,820
@@ 3.3 Prompting and Standards
Given the meme, with the text [$\mathcal{T}$] accompanied by the image [$\mathcal{I}$], is this meme [***]?
This deductive approach in a binary manner is designed to yield more precise and task-specific evaluations from the LMMs, facilitating a clearer understanding of their capability to identify potentially abusive usage of memes. For those still ambiguous outputs, we consider them indicative of the model’s inability to make a nuanced safety judgment regarding the meme. Notably, as different LLMs vary in their adherence to directive outputs, we fine-tune our prompt for each model based on the specific context to achieve controlled outputs. Moreover, we further manually draft rules with regex for filtering, to ensure aligning the model’s responses with one of the two categories in our binary classification standard, thereby facilitating a more structured and effective evaluation process.
@@ 4.2 Main Results
4.3 Effect of CoT Prompts
[…]
Given the meme, with the text [$\mathcal{T}$] accompanied by the image [$\mathcal{I}$], let’s think step by step, is this meme [***]?
Table 3: Results of different LMMs with CoT prompts on the GOAT-Bench across different tasks.
@@ 4.4 Case Study
To better understand the behavior of the best-performed model GPT-4V and facilitate future studies on GOAT-Bench, we provide case studies on both the wrongly and the correctly predicted memes by GPT-4V. To make the analysis more intuitive, we conduct abductive reasoning for each example to obtain the output explanation by GPT-4V for the prediction, as shown in Figure 4-5.

### v4

@@ Abstract
The exponential growth of social media has profoundly transformed how information is created, disseminated, and absorbed, exceeding any precedent in the digital age. Regrettably, this explosion has also spawned a significant increase in the online abuse of memes. Evaluating the negative impact of memes is notably challenging, owing to their often subtle and implicit meanings, which are not directly conveyed through the overt text and image. In light of this, large multimodal models (LMMs) have emerged as a focal point of interest due to their remarkable capabilities in handling diverse multimodal tasks. In response to this development, our paper aims to thoroughly examine the capacity of various LMMs (e.g., GPT-4o) to discern and respond to the nuanced aspects of social abuse manifested in memes. We introduce the comprehensive meme benchmark, GOAT-Bench, comprising over 6K varied memes encapsulating themes such as implicit hate speech, sexism, and cyberbullying, etc. Utilizing GOAT-Bench, we delve into the ability of LMMs to accurately assess hatefulness, misogyny, offensiveness, sarcasm, and harmful content. Our extensive experiments across a range of LMMs reveal that current models still exhibit a deficiency in safety awareness, showing insensitivity to various forms of implicit abuse. We posit that this shortfall represents a critical impediment to the realization of safe artificial intelligence. The GOAT-Bench and accompanying resources are publicly accessible at https://goatlmm.github.io/, contributing to ongoing research in this vital field.
@@ 1 Introduction
Remarkable progress has been made in the field of natural language processing (NLP), marked by the emergence of large language models (LLMs) equipped with billions of parameters and pre-trained on extensive data (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a, b). Through instruction fine-tuning and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), cutting-edge LLMs, like ChatGPT and GPT-4 (OpenAI, 2023), have demonstrated exceptional proficiency in comprehending and executing human instructions, even without tuning the parameters to offer powerful zero-shot or few-shot performance (Wei et al., 2021; Kojima et al., 2022; Lin et al., 2022; Wei et al., 2023). In a related vein, efforts have been directed towards equipping large multimodal models (LMMs) (Team et al., 2023; OpenAI, 2023; Zhu et al., 2023; Liu et al., 2023a; Gong et al., 2023; Dai et al., 2023) with similar capabilities, extending their proficiency to a range of multimodal tasks (Hudson and Manning, 2019; Lu et al., 2022; Fu et al., 2023). This dual advancement in both unimodal and multimodal models marks a significant stride in the field of artificial intelligence, paving the way for more integrated and versatile applications.
@@
Recent literature (Huang et al., 2023) has highlighted that LLMs can achieve remarkable performance in addressing the social abuse inherent in natural languages (An et al., 2021). This area of application is notably more ‘subjective’ compared to other NLP tasks on social media (Lin et al., 2021) and presents significant challenges, even for human evaluators, who often demonstrate lower agreement rates. In addition to language expression, the proliferation of social media has led to the rise of a distinctive multimodal phenomenon: memes (Zenner and Geeraerts, 2018). A typical meme consists of an image coupled with a brief textual component (Pramanick et al., 2021a; Lin et al., 2023). Their ease of sharing enables memes to rapidly disseminate across various online platforms. While commonly perceived with humor, memes are increasingly being co-opted for online abuse, utilizing a strategic interplay of text and image that often relates to current political and socio-cultural events (Sharma et al., 2022; Lin et al., 2024). This evolving use of memes highlights their potential impact on the digital landscape, especially in the context of contemporary discourse. The challenge in detecting social abuse in memes lies in the subtlety of their content, where the underlying meanings are not immediately apparent in the text and images of the memes (Pramanick et al., 2021b). This task demands extensive commonsense knowledge and reasoning skills, built upon a precise interpretation of both the textual and visual elements of memes.
@@
However, in contrast to LLMs, the abilities and constraints of LMMs in handling relatively subjective and nuanced tasks, particularly in identifying online abusive memes within multimodal inputs, are yet to be fully explored (Awadalla et al., 2023; Chen et al., 2023). These tasks are integral to understanding social phenomena, demanding nuanced social judgment and decision-making skills. Given that LMMs are trained on extensive and diverse image-text corpora, showcasing impressive generalization skills (Liu et al., 2023a; Chen et al., 2024), it becomes imperative to assess both the potential strengths and challenges they may face in addressing meme-based social abuse. This line of inquiry is especially pertinent to safety insights, focusing on how LMMs discern and process the intricate combination of visual and textual elements in a manner that is both accurate and socially conscientious so that its output does not cause harm or discomfort.
@@
To this end, we introduce the GOAT benchmark, i.e., GOAT-Bench111 Goat is often depicted as evil in memes due to its historical and mythological associations with devilish figures, reflected in its unique features like horns and eyes., a comprehensive and specialized dataset designed to evaluate large multimodal models through meme-based multimodal social abuse. GOAT-Bench comprises over 6K diverse memes, encompassing a range of themes including hate speech and offensive content. Our focus is to assess the ability of LMMs to accurately identify online abuse, specifically in terms of hatefulness, misogyny, offensiveness, sarcasm, and harmfulness. We meticulously control for the granularity of each specific meme task to facilitate a detailed analysis. Furthermore, we extend our evaluation to assess the effectiveness of thought chains, few-shot demonstrations, and supervised fine-tuning in discerning the underlying implications of memes for deducing their potential threat to online safety.
@@
Leveraging the GOAT-Bench, we assess the capability of various LMMs as exemplified in Figure 1, to detect meme-based social abuse through complex reasoning. We conceive of GOAT-Bench as a comprehensive suite of benchmarks, with subsets tailored to evaluate specific model abilities and analyze the key strengths and weaknesses of LMMs. For instance, in tasks focused on sexism, such as misogyny, we observe significant disparities among LMMs; in broader tasks like harmfulness, which demand advanced background knowledge and reasoning, the performance of LMMs tends to be more concentrated and generally modest. Our aim with GOAT-Bench is to enable researchers to gain a multi-dimensional understanding of their models’ abilities in addressing meme-based social abuse and to contribute to the advancement of safety insights in LMMs, as a critical step in preventing the escalation of online social abuse and ensuring the stability and cohesion of diverse communities.
@@
Our contributions are summarized as follows in three folds:
  • We introduce GOAT-Bench, the first advanced comprehensive testbed consisting of 6,626 memes, spanning five tasks of varied complexity related to social abuse, tailored to systematically dissect and analyze the safety insights of LMMs.
  • We extensively evaluate the performance of the 11 most cutting-edge large multimodal models on the GOAT-Bench benchmark. Our findings indicate that GOAT-Bench poses significant challenges to the current state-of-the-art models. Remarkably, the top-performing GPT-4V only attains an overall macro-averaged F1 score of 70.29%, with all the other open-source LMMs registering scores below 62%.
  • Our study conducts an empirical and exhaustive analysis of the performance disparities among various closed-source and open-source LMMs on GOAT-Bench, employing diverse deployment strategies such as chain-of-thought, in-context learning prompts, and the instruction-tuning paradigm. It is our aspiration that our research, along with the GOAT-Bench benchmark, lays the groundwork for future advancements in this realm, catalyzing the evolution of increasingly sophisticated and proficient large multimodal models.
@@ 2.1 Tasks
Task III: Offensiveness. Offensiveness spreads an idea or emotion that intends to damage the social identity of the target person, community, or lower their prestige (Suryawanshi et al., 2020a). In contrast to hateful memes, which target specific groups with expressions of animosity or discrimination, offensive memes encompass a wider spectrum of language or image that has the potential to offend, hurt, or provoke anger among individuals. They may not be directed towards a particular group and can affect individuals in a general sense (Drakett et al., 2018). In light of this understanding, it becomes imperative to investigate the efficacy of LMMs in scrutinizing online multimodal offensive meme content, known for its intent to provoke distress or embarrassment through its use of rudeness. To this end, we organized the offensive memes by utilizing the MultiOFF dataset (Suryawanshi et al., 2020a) related to the 2016 United States presidential election, which consists of abusive and discriminatory messages against a person or minority group.
@@
Term Scope Sources #Distribution of Labels #Total
@@
Misogyny Target women with sexism or hate, involving shaming, stereotyping, objectification, or violence. MAMI Misogynistic 500 1,000
@@
Sarcasm Feature irony and humor to subtly express dissatisfaction or mock, often through incongruous text and image. MSD Sarcastic 910 1,820
@@ 3.3 Prompting and Standards
Given the meme, with the text [$\mathcal{T}$] accompanied by the image [$\mathcal{I}$], is this meme [*]?
This deductive approach in a binary manner is designed to yield more precise and task-specific evaluations from the LMMs, facilitating a clearer understanding of their capability to identify potentially abusive usage of memes. For those still ambiguous outputs, we consider them indicative of the model’s inability to make a nuanced safety judgment regarding the meme. Notably, as different LLMs vary in their adherence to directive outputs, we adapt our prompt for each model based on the specific context to achieve controlled outputs. Moreover, we further manually draft rules with regex for filtering, to ensure aligning the model’s responses with one of the two categories in our binary classification standard, thereby facilitating a more structured and effective evaluation process.
@@ 4.2 Main Results
4.3 Effect of Zero-Shot CoT Prompts
Table 3: Results of different LMMs with zero-shot CoT prompts on the GOAT-Bench across different tasks.
@@ 4.3 Effect of Zero-Shot CoT Prompts
Model (#Params) Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$
@@
[…]
Given the meme, with the text [$\mathcal{T}$] accompanied by the image [$\mathcal{I}$], let’s think step by step, is this meme [*]?
@@
4.4 Effect of Few-Shot In-context Learning
[…]
You are a detector trained to identify meme-based social abuse. Here are some examples: **example1** Given a meme: [example1], is this meme [*]? Answer: Yes. **example2** Given a meme: [example2], is this meme [*]? Answer: No. Based on the meme examples, now given the meme, with the text [$\mathcal{T}$] accompanied by the image [$\mathcal{I}$], is this meme [*]?
It is noteworthy that 1) the LMM’s performance significantly declines with an increase in the number of few-shot, in-context learning examples. This decline may be attributed to the LMMs’ difficulty in abstracting generalized reasoning patterns from the complex, multimodal thinking processes inherent in a limited set of meme examples. This suggests that while the LLaVA can handle straightforward tasks with minimal data, their ability to scale up reasoning capabilities in more intricate scenarios is limited. 2) Simultaneously, except for the other four meme tasks, LLaVA only manages to glean some performance improvements in the finer-grained Misogyny task from a finite set of ICL demonstrations, yet this improvement is marked by a degree of fluctuation due to the unpredictable nature of meme content and context. This indicates that while LLaVA can adapt to specific, narrowly defined tasks, its performance is not consistently reliable across broader or more variable contexts. 3) Overall, the in-context learning prompting strategy does not yield a consistent upward trend in performance or enhance the model’s robustness. This reflects the formidable challenge meme-based social abuse poses to the LMM paradigm, underscoring the intricate complexities and nuanced understandings required to effectively navigate and interpret such content in future multimodal research about large language models. We speculate that the primary reason for these challenges is likely due to the model’s lack of sufficient training on multi-image scenarios. This insufficiency means that LMMs are not fully equipped to handle the diverse and layered contexts presented by memes, which often combine text, images, and cultural references in ways that are difficult to parse without extensive, specialized training. Addressing this gap will be crucial for future developments in multimodal learning, as it highlights the need for training data that better reflects the complexity of real-world applications.
Table 4: Results of LLaVA with ICL prompts on the GOAT-Bench across different tasks.
#Examples Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$ Acc. $\emph{F}_{1}$
0-shot 65.20 61.40 57.10 48.68 58.68 58.39 73.52 72.89 66.70 64.35 65.77 61.14
1-shot 62.70 60.02 68.70 68.05 46.30 40.50 63.08 62.98 47.47 44.52 59.43 55.21
2-shot 63.00 42.19 58.40 52.74 59.49 52.44 62.47 60.72 56.29 56.27 60.69 52.87
3-shot 62.80 40.06 55.20 45.71 60.97 48.37 58.57 52.39 59.56 55.38 59.77 48.38
4-shot 63.20 50.06 67.60 66.62 49.66 48.05 52.69 44.68 53.02 52.49 57.83 52.38
5-shot 63.00 55.96 69.10 69.08 46.97 42.52 51.92 44.57 53.22 52.40 57.51 52.91
6-shot 40.60 35.11 52.10 51.31 39.57 35.89 48.30 39.23 41.33 40.71 44.45 40.45
@@ 4.5 Case Study
To better understand the behavior of the best-performed model GPT-4V and facilitate future studies on GOAT-Bench, we provide case studies on both the wrongly and the correctly predicted memes by GPT-4V. To make the analysis more intuitive, we conduct abductive reasoning for each example to obtain the output explanation by GPT-4V for the prediction, as shown in Figure 4-6.
@@
Figure 5: Non-hateful example of wrongly predicted memes by GPT-4V with the explanation.